OpenStudy (anonymous):

Use inductive reasoning to find the next term in the sequence 2, 3, 5, 8, 12 … What is the next term? A. 13 B. 15 C. 17 D. 18 @OrangeMaster

OpenStudy (anonymous):

c.17

OpenStudy (anonymous):

17

OpenStudy (anonymous):

2+1=3 3+2=5 5+3=8 8+4=12 12+5=17